# Random time-shift approximation enables hierarchical Bayesian inference of mechanistic within-host viral dynamics models on large datasets

### v2

@@ 3.1 Simulation study
Table 4 reports numerical summaries of the posterior fits, while Fig. 3 shows the 95% credible intervals for the population parameters. The true values fall within these intervals in most cases, with the lowest empirical coverage at 0.90. The credible intervals are narrow relative to the scale of the parameters (Interval width column in Table 4), indicating that the model recovers the population values with reasonable precision. Relative bias was computed as the bias divided by the true parameter value, giving a signed percentage difference. For all parameters except $\sigma_{R_{0}}$ and $\sigma_{\rho}$, posterior means are within $\pm 3.7\%$ of the true value. For $\sigma_{R_{0}}$ and $\sigma_{\rho}$ the relative biases are 0.114-0.114 and $0.626$, respectively. However, Fig. 3 shows that the 95% credible intervals for these variance parameters are much wider (average widths 1.036 and 0.827), reflecting substantial uncertainty. This explains why, despite noticeable bias in the posterior means of these two parameters, interval coverage remains adequate.
@@
$\mu_{R_{0}}$ $0.90$ 0.964[0.901,1.051]0.964\>[0.901,1.051] 0.021[0.013,0.059]0.021\>[-0.013,0.059]
$\sigma_{R_{0}}$ $1.00$ 1.036[0.808,1.233]1.036\>[0.808,1.233] 0.114[0.397,0.23]-0.114\>[-0.397,0.23]
$\mu_{\delta}$ $0.920$ 0.086[0.081,0.092]0.086\>[0.081,0.092] 0.008[0.027,0.013]-0.008\>[-0.027,0.013]
$\sigma_{\delta}$ $0.90$ 0.056[0.052,0.06]0.056\>[0.052,0.06] 0.037[0.162,0.1]-0.037\>[-0.162,0.1]
$\mu_{\rho}$ $0.920$ 0.750[0.685,0.823]0.750\>[0.685,0.823] 0.017[0.053,0.083]0.017\>[-0.053,0.083]
$\sigma_{\rho}$ $0.980$ 0.827[0.665,1.015]0.827\>[0.665,1.015] 0.626[0.011,1.643]0.626\>[-0.011,1.643]
$\kappa$ $0.940$ 0.044[0.042,0.045]0.044\>[0.042,0.045] 0.006[0.025,0.027]0.006\>[-0.025,0.027]
Table 4: Numerical summary of posterior fits for the shared parameters. Coverage is the proportion of the 50 simulations in which the true parameter value lies within the 95% credible interval. The interval width is averaged across the 50 simulations, with an 80% confidence interval derived from quantiles ($q_{10},q_{90}$) reported. Relative bias is computed by dividing the bias by the true value to capture the scale of the difference, again with 80% confidence interval provided. An 80% confidence level was chosen to provide adequate precision given only 50 observations (datasets).
Fig 3: 95% credible interval plots of the population parameters for the simulated study. Blue lines indicate the 95% credible interval for a given simulation, with blue points indicating the median. The horizontal red dashed line shows the true value of the parameters used throughout the simulations.
@@ 3.2 NBA study
The marginal posteriors for the shared parameters are shown in Fig. 6. We see that there is large movement in the distributions from the prior (orange) to the posteriors (blue) for all parameters besides $\sigma_{R_{0}}$. This is not surprising since the individual parameters $R_{0}$, $\delta$, and $\rho$ values trade-off with one another and with less data in the growth phase, this can cause the related hyper-parameters to be weakly identifiable when inferred jointly. The model appears to be able to estimate all parameters including the measurement noise $\kappa$. Our estimates for $\mu_{R_{0}}$—and by extension—the individual $R_{0}$’s are consistent with the results in Table 2. It should be noted that compared to Zitzmann et al. [5, SI] our population estimate is larger, $13.4$ compared with $8.2$. Our $R_{0}$ estimates are consistent with the value of 14.2 reported in Germano et al. [18] from their stochastic model (which also accounts for immune responses). Our population estimate for $\mu_{\rho}$ is $3.27$ which is slightly larger than the estimate of $3.07$, but still consistent with [5, SI]. We infer the population mean rate of removal for infectious cells to be $1.18$, slightly lower than the $1.3$ [5, SI]. There are several possible reasons for the difference in the estimates of $R_{0}$ (i.e. $\mu_{R_{0}}$). The first being that, as our model is stochastic, larger values of the parameters are required to ensure that the model overcomes extinction than what would be estimated by a deterministic model. We are also fitting to a much larger number of individuals (163 vs 25) including those that are unvaccinated and have limited data during the growth and decline phases. Additionally, we account for the LOD in our model and make no assumptions as to the time from infection to peak instead allowing our model to infer the infection time.
Fig 6: Density plots for the marginal distributions of the shared parameters (hyper-parameters and the scale of the measurement noise) for the NBA model fit. Prior distributions are indicated in orange and posteriors are given by the blue histograms.
@@ 4 Discussion
We have presented a novel method for fitting mechanistic within-host viral dynamics models that explicitly account for process noise, providing a more realistic representation of viral load (VL) trajectories. Our approach leverages the concept of a random time-shift—first introduced in Barbour et al. [16] and later developed into a practical tool in Morris et al. [17]—to capture the effects of stochasticity in the WHVDs. Our key contribution in this work is demonstrating how the time-shift distributions can be directly incorporated into the likelihood function, allowing stochasticity to be modelled without the need for computationally expensive stochastic simulations, for example using Gillespie’s algorithm [22] or tau-leaping [23]. Furthermore, this enables efficient inference, making it feasible to fit the model to large datasets using a hierarchical framework. Our approach can be viewed as a hybrid method, combining the interpretability and tractability of deterministic models while capturing key elements of stochastic modelling, offering a middle ground between fully stochastic simulation and purely deterministic fitting.
@@
Hierarchical approaches have been widely used in phenomenological studies of WHVDs [6, 11], but to our knowledge, no stochastic WHVD models have incorporated hierarchical structures. The existing approaches fall into two main categories: simpler parametric models, such as piecewise linear functions [4, 11, 6], and mechanistic models that assume deterministic dynamics [12, 5, 40, 25]. Parametric models offer computational advantages, as their likelihoods are easier to evaluate and they integrate well with probabilistic programming frameworks like Stan [55]. However, they primarily describe viral dynamics in terms of slope parameters (for growth and decline) and a hinge point (for peak VL), limiting their ability to extract biologically meaningful quantities like $R_{0}$ and production rates $\rho$. Mechanistic models, in contrast, provide a more realistic representation of infection dynamics but are computationally expensive due to the need to repeatedly solve ODEs. This challenge is especially pronounced in Bayesian inference for large datasets. Moreover, parameter identifiability can become a major concern in this context, particularly when biological parameters influence the VL in overlapping ways (i.e. the relationship between $R_{0}$, the viral production rate, the infection time and initial conditions) [5, 25]. While some hybrid stochastic-deterministic models [56, 57, 18] have been developed, they remain computationally demanding and have not been widely applied to multi-individual VL kinetics studies. Our method is the first to use a mechanistic model within a hierarchical framework that can also scale to incorporate hundreds of individuals.
@@
We have adopted the simplest viral dynamics model for this work due to parameter identifiability concerns, but the method can easily be applied to other suitable models. These models must satisfy the conditions required for applying the time-shift approximation [17, 16], which many biological systems with large populations naturally fulfill [58, 17, 59]. The main constraint beyond the large population assumption lies in data availability during the pre-exponential growth phase. Systems where early-time dynamics are critical to macroscopic outcomes but where direct observations are limited—such as other within-host models—are particularly well suited for this approach. Examples include the innate response model used in Zitzmann et al. [5] and more complex models, such as those examining the role of MUC1 in reducing influenza severity [14]. Conversely, these conditions also define when our method may not be the most appropriate. In cases where high-quality early-growth-phase data is available, exact stochastic methods may be preferable. Epidemic outbreak data, for instance, typically includes detailed observations during the initial growth phase when cases are few and monitoring efforts are robust, making alternative approaches more suitable.

### v3

@@ 3.1 Simulation study
Table 4 reports numerical summaries of the posterior fits (see Section 7 of S1 text. for details on each of the scoring rules shown in Table 4). Figure 3 displays the 95% credible intervals for the population parameters across 50 of the 200 simulated datasets. In most cases, the true values are contained within the 95% credible intervals, with the lowest empirical coverage being 0.95. The credible intervals are narrow relative to the scale of the parameters (see the Interval width column in Table 4), indicating that the model recovers the population values with good precision. Relative bias was computed as the bias divided by the true parameter value, giving a signed percentage difference. For all parameters, zero lies within the 95% confidence interval for the relative bias, suggesting that the method is unbiased. For $\sigma_{R_{0}}$ and $\sigma_{\rho}$, the average relative biases are 0.109-0.109 and $0.585$, respectively. However, Figure 3 shows that the 95% credible intervals for these variance parameters are much wider (average interval widths of 1.05 and 0.826, respectively), reflecting substantial uncertainty. This explains why, despite noticeable bias in the posterior means of these two parameters, interval coverage remains adequate.
@@
$\mu_{R_{0}}$ $0.95$ 0.958[0.862,1.096]0.958\>[0.862,1.096] 0.016[0.032,0.072]0.016\>[-0.032,0.072]
$\sigma_{R_{0}}$ $1.00$ 1.05[0.808,1.233]1.05\>[0.808,1.233] 0.109[0.422,0.525]-0.109\>[-0.422,0.525]
$\mu_{\delta}$ $0.95$ 0.086[0.078,0.095]0.086\>[0.078,0.095] 0.007[0.037,0.025]-0.007\>[-0.037,0.025]
$\sigma_{\delta}$ $0.95$ 0.056[0.05,0.063]0.056\>[0.05,0.063] 0.021[0.215,0.146]-0.021\>[-0.215,0.146]
$\mu_{\rho}$ $0.97$ 0.745[0.654,0.858]0.745\>[0.654,0.858] 0.020[0.093,0.134]0.020\>[-0.093,0.134]
$\sigma_{\rho}$ $0.98$ 0.826[0.774,1.443]0.826\>[0.774,1.443] 0.585[0.106,2.067]0.585\>[-0.106,2.067]
$\kappa$ $0.96$ 0.044[0.041,0.046]0.044\>[0.041,0.046] 0.006[0.036,0.044]0.006\>[-0.036,0.044]
Table 4: Numerical summary of 200 posterior fits for the shared parameters. Coverage is the proportion of simulations (out of 200) in which the true parameter value lies within the 95% credible interval. The average interval width is reported across simulations, with a 95% confidence interval. Relative bias is also reported with its corresponding 95% confidence interval. See Section 7 of S1 text. for further details of the scoring metrics.
Fig 3: 95% credible interval plots of the population parameters for 50 of the 200 simulated studies. Blue lines indicate the 95% credible interval for a given simulation, with blue points indicating the median. The horizontal red dashed line shows the true value of the parameters used throughout the simulations.
@@ 3.2 NBA study
The marginal posteriors for the shared parameters are shown in Fig. 6. We see that there is large movement in the distributions from the prior (orange) to the posteriors (blue) for all parameters. The model appears to be able to estimate all parameters including the measurement noise $\kappa$. Posterior means and 95% credible intervals are provided in Table 5. Our estimates for $\mu_{R_{0}}$ are consistent with the results in Table 2. It should be noted that compared to Zitzmann et al. [5, SI] our population estimate is larger, $14.64$ compared with $8.2$. Our $R_{0}$ estimates are consistent with the value of 14.2 reported in Germano et al. [18] from their stochastic model (which also accounts for immune responses). Our population estimate for $\mu_{\rho}$ is $3.19$ which is slightly larger than, but still consistent with the estimate of $3.07$ from [5, SI]. We infer the population mean rate of removal for infectious cells to be $1.15$, slightly lower than $1.3$ in [5, SI]. There are several possible reasons for the difference in the estimates of $R_{0}$ (i.e. $\mu_{R_{0}}$). The first being that, as our model is stochastic, larger values of the parameters are required to ensure that the model overcomes extinction than what would be estimated by a deterministic model. We are also fitting to a much larger number of individuals (163 vs 25) including those that are unvaccinated and have limited data during the growth and decline phases. Additionally, we account for the LOD in our model and make no assumptions as to the time from infection to peak instead allowing our model to infer the infection time.
Parameter Mean (95% CrI)
$\mu_{R_{0}}$ $14.64$ [12.23, 17.60]
$\sigma_{R_{0}}$ $4.34$ [3.07, 5.98]
$\mu_{\delta}$ $1.15$ [1.07, 1.23]
$\sigma_{\delta}$ $0.19$ [0.12, 0.26]
$\mu_{\rho}$ $3.19$ [2.28, 4.35]
$\sigma_{\rho}$ $1.35$ [0.71, 2.10]
$\kappa$ $1.29$ [1.23, 1.37]
Table 5: Posterior means and 95% credible intervals for the shared parameters.
Fig 6: Histograms of the marginal posterior distributions for the shared parameters (hyper-parameters and the scale of the measurement noise) for the NBA model fit. Prior distributions are indicated by orange lines and posteriors are shown as blue histograms.
@@ 4 Discussion
We have presented a novel method for fitting mechanistic within-host viral dynamics models that explicitly account for process noise, providing a more realistic representation of viral load (VL) trajectories. Our approach leverages the concept of a random time-shift—first introduced in Barbour et al. [16] and later developed into a practical tool in Morris et al. [17]—to capture the effects of stochasticity in the WHVDs. Our key contribution in this work is demonstrating how the time-shift distributions can be directly incorporated into the likelihood function, allowing stochasticity to be modelled without the need for computationally expensive stochastic simulations, for example using Gillespie’s algorithm [22] or tau-leaping [23]. Furthermore, this enables efficient inference, making it feasible to fit the model to large datasets using a hierarchical framework. Hierarchical approaches have been widely used in phenomenological studies of within-host viral dynamics (WHVDs) [6, 11], but to our knowledge, no stochastic WHVD models have incorporated hierarchical structures. Our approach can be viewed as a hybrid method, combining the interpretability and tractability of deterministic models while capturing key elements of stochastic modelling, offering a middle ground between fully stochastic simulation and purely deterministic fitting.
@@
Existing approaches for fitting WHVDs fall into two main categories: simpler phenomenological models, such as piecewise linear functions [4, 11, 6], and mechanistic models that assume deterministic dynamics [12, 5, 40, 25, 46, 44]. Phenomenological models offer computational advantages, as their likelihoods are easier to evaluate and they integrate well with probabilistic programming frameworks like Stan [56]. However, they primarily describe viral dynamics in terms of slope parameters (for growth and decline) and a hinge point (for peak VL), limiting their ability to infer biologically meaningful quantities such as the basic reproduction number $R_{0}$ or production rate $\rho$. Mechanistic models, by contrast, provide a more realistic representation of infection dynamics but are computationally more demanding, as they require repeatedly solving ODEs or simulating stochastic processes. Such ODE-based models can be fit using dedicated mixed-effects tools such as Monolix [54], which implement efficient frequentist approaches for nonlinear hierarchical models but are less flexible than Bayesian frameworks for incorporating stochastic dynamics or propagating uncertainty in a principled way. Parameter identifiability also poses a major challenge, particularly when biological parameters influence viral load in overlapping ways (e.g. relationships among $R_{0}$, the viral production rate, infection timing, and initial conditions) [5, 25]. Although some hybrid stochastic–deterministic models have been developed [57, 58, 18], they remain computationally intensive and have not been widely applied to large, multi-individual viral kinetics datasets. Our method is the first to use a stochastic mechanistic model within a hierarchical framework that scales efficiently to datasets with over a hundred individuals.
@@
We have adopted the simplest viral dynamics model for this work due to parameter identifiability concerns, but the method can easily be applied to other suitable models. These models must satisfy the conditions required for applying the time-shift approximation [17, 16], which many biological systems with large populations naturally fulfill [59, 17, 60]. The main constraint beyond the large population assumption lies in data availability during the pre-exponential growth phase. Systems where early-time dynamics are critical to macroscopic outcomes but where direct observations are limited, such as other within-host models, are particularly well suited for this approach. Examples include the innate response model used in Zitzmann et al. [5] and more complex models, such as those examining the role of MUC1 in reducing influenza severity [14]. Conversely, these conditions also define when our method may not be the most appropriate. In cases where high-quality early-growth-phase data is available, exact stochastic methods may be preferable. Epidemic outbreak data, for instance, typically includes detailed observations during the initial growth phase when cases are few and monitoring efforts are robust, making alternative approaches more suitable.